I am trying to help my granddaughter and need to explane this to her.
x + 4 = y -2 y = x + x + x

I got x = 2 and y = 6 but was told it was wrong. Help Please

Answer by checkley71(8403)

Y=3X & X+4=Y-2 SO X+4=3X-2 OR X-3X=-2-4 OR -2X=-6 OR 2X=6 OR X=3 THEN Y=3*3 OR
Y=9 THEN 3+4=9-2 OR 7=7